Find l^-1 {1/(s+1)(s+2)^2 by using convolution theorem
1
Expert's answer
2020-09-21T15:02:31-0400

By Convolution Theorem :


"\\
L^{-1}\{F_{1}(s).F_{2}(s)\}=\int_0^tF_{1}(u).F_{2}(t-u) \;du"


Here,we need to evaluate "L^{-1}\{\dfrac{1}{(s+1)(s+2)^2}\}=L^{-1}\{\dfrac{1}{s+1}.\dfrac{1}{(s+2)^2}\}" :


Step-1 : Take "F_1(s)=\dfrac{1}{(s+1)} \;and\; F_2(s)=\dfrac{1}{(s+2)^2}"


Step-2 : "L^{-1}\{F_1(s)\}=L^{-1}\{\dfrac{1}{(s+1)}\}"

Using the properties of inverse laplace transform : "L^{-1}\{\dfrac{1}{s-a}\}=e^{at}\;\;and\;\;L^{-1}\{\dfrac{1}{s^n}\}=\dfrac{t^{n-1}}{(n-1)!}"


"\therefore L^{-1}\{F_1(s)\}= L^{-1}\{\dfrac{1}{(s+1)}\}=e^{-1.t}=e^{-t}=\mathbf{f_1(t)}\;and"



"L^{-1}\{F_2(s)\}=L^{-1}\{\dfrac{1}{(s+2)^2}\}=e^{-2t}L^{-1}\{\dfrac{1}{s^2}\}"


"\implies L^{-1}\{F_2(s)\}=e^{-2t}.\dfrac{t^{2-1}}{(2-1)!}=e^{-2t}.\dfrac{t}{1!}=te^{-2t}=\mathbf{f_2(t)}"


By using the convolution formula,we have -


"L^{-1}\{\dfrac{1}{(s+1)(s+2)^2}\}=L^{-1}\{\dfrac{1}{s+1}.\dfrac{1}{(s+2)^2}\}"


"=\int_0^t\mathbf{f_1(u).f_2{(t-u)du}}"


"=\int_0^te^{-u}.(t-u).e^{-2(t-u)}du"


"=\int_0^t(t-u).e^{-u-2t+2u}du"


"=\int_0^t(t-u).e^{u-2t}du=e^{-2t}\int_0^t(t-u)e^udu"


"=e^{-2t}[t.\int_0^te^udu-\int_0^tue^udu]"


"=e^{-2t}\bigg[t.\bigg[e^u\bigg]_0^t-\bigg[u.\int e^udu-\int \bigg(\dfrac{d}{du}(u).\int e^udu\bigg)du\bigg]_0^t\bigg]"


"=e^{-2t}\bigg[t.[e^t-e^0]-\bigg[u.e^u-\int\bigg(1.e^u\bigg)du\bigg]_0^t\bigg]"


"=e^{-2t}\bigg[t.[e^t-1]-\bigg[ue^u-e^u\bigg]_0^t\bigg]"


"=e^{-2t}\bigg[te^t-t-\bigg[(te^t-e^t)-(0-e^0)\bigg]\bigg]"


"=e^{-2t}\bigg[te^t-t-\bigg[(te^t-e^t)+1)\bigg]\bigg]"


"=e^{-2t}\bigg[te^t-t-te^t+e^t-1\bigg]=e^{-2t}(e^t-t-1)" .............Ans.